# Autoregressive Moving-average Attention Mechanism for Time Series Forecasting

### v1

@@ 3 Experiments
Table 2: Summary of main TSF results with forecasting horizons $L_{P}\in\{12,24,48,96\}$ and $L_{I}=512$. See Table 8 for the original results. Averages of test set MSE for each model on each dataset are presented. Average rankings (AvgRank) of each model, along with the count of first-place rankings (#Top1), are also included. Green indicates better performance, while red indicates worse. For each comparison between AR and ARMA attention, the better model is underlined.
Model AR/ARMA Transformer Baseline
@@
Table 3: Summary showing that ARMA Transformers with $m=3$ layers consistently outperform their AR counterparts across a wide range of $m$. The same experimental settings and data presentation method as in Table 2 are used. See Table 9 for the original results.
@@
      ARMA       AR       AR       AR       AR       AR       AR       AR       AR
@@
Performance and complexity   The improvement of adding the MA term comes from its ability to model short-term impacts, allowing the AR term to focus on long-term and cyclic effects, not from increased computational costs. Table 3 shows that, regardless of the number of layers $m$ (1 to 8), AR Transformers consistently underperform compared to ARMA Transformers with a fixed $m=3$.
Table 4: Summary showing that AR/ARMA Transformers effectively utilize extended lookback $L_{I}$, while baselines experience performance degradation. $L_{I}\in\{512,1024,2048,4096\}$ with $L_{P}\in\{12,24,48,96\}$ are evaluated and averaged. Original results can be found in Table 10.
Model AR/ARMA Transformer Baseline
@@
Adaptability to Longer $L_{I}$   Previous baseline models typically use $L_{I}$ between 96 and 720, as longer $L_{I}$ often leads to overfitting to long-term patterns, ignoring more important local effects (Zeng et al., 2023; Nie et al., 2022; Liu et al., 2024a). However, the next-step prediction and varying lookback inputs in AR/ARMA Transformers help the model focus on tokens closer to the next step, improving generalization. As shown in Table 4, increasing $L_{I}$ from 512 to 4096 improves AR/ARMA performance, demonstrating scalability and the ability to properly leverage long-term effects. Also, the ARMA structure consistently boosts AR model performance across different $L_{I}$.
@@
Varying $L_{P}$   We evaluated AR/ARMA Transformers with $L_{I}=4096$ on long-term output $L_{P}=720$ while keeping the best-performing $L_{I}=512$ for the baselines. We also compared short-term output with $L_{P}=1$ using $L_{I}=512$ for all models. As shown in Table 5, across different $L_{P}$, AR models performed comparably to baselines, and the ARMA structure consistently improved performance over AR models.
Table 5: Summary of model performance on varying horizons $L_{P}$. AR/ARMA Transformers uses $L_{I}=512$ for $L_{P}=1$ and $L_{I}=4096$ for $L_{P}=720$. Baselines are consistently set to their best-performing $L_{I}=512$ configuration. Original results can be found in Table 11.
Model AR/ARMA Transformer Baseline
+ARMA +ARMA +ARMA +ARMA +ARMA former Former
Weather $L_{P}=1$ 0.037 0.032 0.032 0.031 0.033 0.031 0.037 0.034 0.036 0.034 0.038 0.038 0.033 0.035 0.034 0.032
$L_{P}=720$ 0.299 0.301 0.308 0.305 0.310 0.299 0.296 0.295 0.299 0.297 0.327 0.317 0.311 0.325 0.319 0.404
ETTm1 $L_{P}=1$ 0.048 0.043 0.042 0.041 0.051 0.043 0.053 0.051 0.054 0.052 0.047 0.049 0.043 0.046 0.044 0.044
$L_{P}=720$ 0.396 0.391 0.417 0.408 0.396 0.393 0.395 0.391 0.403 0.399 0.420 0.438 0.418 0.408 0.420 0.802
ETTm2 $L_{P}=1$ 0.034 0.031 0.032 0.030 0.033 0.030 0.035 0.033 0.033 0.031 0.034 0.034 0.032 0.033 0.031 0.035
$L_{P}=720$ 0.329 0.327 0.341 0.338 0.334 0.328 0.327 0.326 0.331 0.327 0.360 0.369 0.351 0.362 0.403 2.641
@@
Visualization analysis   Fig. 5 shows test loss curves for different AR/ARMA attention mechanisms on the Weather and ETTm1 datasets, with ARMA consistently outperforming AR in both convergence speed and final loss. Fig. 6 visualizes attention input sequence, AR weights, $\mathbf{B}$, and $\mathbf{\Theta}$ matrices of a test datapoint on Weather, showing how MA weights decouple local patterns, allowing AR weights to focus on cyclic and long-term patterns. Additional visualizations in Figs. 912 reinforce that there are important long-term stable seasonal patterns for AR weights to capture that should not be disrupted by applying forget gates or EMA. This explains why gated linear attention underperforms linear attention in our experiments.
Figure 5: Visualization of test loss curves. We show the testing performance of five attention mechanisms using AR/ARMA structures on the Weather and ETTm1 datasets ($L_{I}=512$, $L_{P}=48$).
Figure 6: Visualization of the ARMA attention weights (first attention layer, averaged across the multiple heads or $d$-dimensional channels) for the first test set data point in the Weather dataset ($L_{I}=4096$, $L_{P}=96$). More ARMA weight visualization can be found in Fig. 9, 10, 11, and 12.
@@
Computational cost   Table 7 compares the computational cost of AR/ARMA Transformers with baselines on the ETTm1 dataset. Our tokenization method reduces the token size $N$, keeping AR/ARMA models’ computational cost comparable to the baselines. Additionally, parameter sharing ensures the MA term doesn’t increase the number of parameters, and the extra FLOPs from using ARMA are not significant.
@@
4 Related works
5 Conclusion, limitation, and future works
We propose the ARMA attention mechanism, which integrates an MA term into existing AR attention using a novel indirect MA weight generation method. This approach maintains the same time complexity and parameter size while ensuring the validity of the implicit MA weights. Experiments demonstrate that ARMA attention successfully decouples and handles long-term and short-term effects. The ARMA Transformer, enhanced with the MA term, outperforms their AR counterparts and achieves state-of-the-art results, offering consistent improvements in training with minimal added computational cost.
One limitation is that we have not explored combining the channel-independent ARMA Transformer with multivariate forecasting models to improve its handling of inter-series relationships. For future work, ARMA attention could be applied to general sequence modeling tasks beyond TSF. Testing on larger-scale datasets, such as using ARMA Transformers for large-scale NLP pretraining, is another promising direction.
Appendix A Appendix
A.1 Datasets
Traffic Dataset [^6] http://pems.dot.ca.gov/(Wu et al., 2021)
PEMS Dataset [^7] http://pems.dot.ca.gov/(Li et al., 2017)
A.2 Hyper-parameter settings and implementation details
For the hyper-parameter settings of the AR/ARMA Transformer, we use $m=3$ Transformer layers, 8 heads, and set the hidden dimension $d$ based on the number of series $C$, using the empirical formula $d=16\sqrt{C}$. We use $4d$ as the hidden dimension for the feedforward MLP in the Transformer layer. A dropout rate of 0.1 is applied to both the AR term and MA term. We initialize the weights of all linear layers and embedding layers using the GPT-2 weight initialization method, with a normal distribution and a standard deviation of 0.02. For the output projection layers in the attention and MLP, we additionally scale the standard deviation by a factor of $1/\sqrt{m}$, aligned with the GPT-2 setting. Normalization layer is applied both before the input to the Transformer and after the Transformer output. We experimented with both standard LayerNorm and RMSNorm as the normalization layer, finding no significant performance differences, so we opted for RMSNorm for lower computational cost. For token input projection, we use a linear layer to project the $L_{P}$-dimensional token to a $d$-dimensional input vector. In the output projection, we do not tie the weights between the input and output linear layers. A learnable position embedding that maps the integer labels from 1 to $N$ (the input sequence length) to the corresponding $d$-dimensional position vectors is used. At the beginning of the model, we apply RevIN to input series $\mathbf{S}_{I}$, subtracting the mean and dividing by the standard deviation for each series. Before outputting the final result, we multiply by the standard deviation and add the mean back. All input series are processed independently and in parallel, merging different series dimensions into the batch size for parallel computation. The random seed used in all the experiments is 2024.
All training tasks in this paper can be conducted using a single Nvidia RTX 4090 GPU. The batch size is set to 32. For larger datasets, such as Traffic and PEMS07, we use a batch size of 16 or 8, with 2-step or 4-step gradient accumulation to ensure the effective batch size for parameter updates remains 32. During training, AR/ARMA Transformers are trained using the next-step prediction objective with MSE loss. We use the AdamW optimizer with betas=(0.9, 0.95) and weight decay=0.1, following the GPT-2 settings. For a fair comparison, the same optimizer is used for training baseline models. It is important to note that the baseline models trained with this AdamW setup show significantly better TSF performance compared to those trained with the default Adam optimizer settings. As a result, the baseline performance presented in this paper may exceed the results reported in their original papers. Since this study focuses on long-term last token prediction results, we apply an additional weight factor to the training loss for the last token, multiplying it by $N$. However, this weighting only slightly affects performance on smaller datasets with fewer data points, such as ETTs, and has little to no effect on larger datasets. Given the minimal impact of this method, the original next-token MSE loss is sufficient for most datasets, without requiring further modifications.
A.3 Supplementary experiment results
@@
Table 9: Results showing that ARMA Transformers with $m=3$ layers consistently outperform their AR counterparts across a wide range of $m$. Forecasting horizons $L_{P}\in\{12,24,48,96\}$ and $L_{I}=512$ are used. Test set MSE and MAE for each model on each experiment setup are presented.
Model ARMA(m=3) AR(m=1) AR(m=2) AR(m=3) AR(m=4) AR(m=5) AR(m=6) AR(m=7) AR(m=8)
@@
Table 10: Results showing that AR/ARMA Transformers effectively utilize extended lookback $L_{I}$, while baselines experience performance degradation. $L_{I}\in\{512,1024,2048,4096\}$ with $L_{P}\in\{12,24,48,96\}$ are evaluated. Test set MSE and MAE for each model on each setup are presented.
@@
Table 11: Results of model performance on varying horizons $L_{P}$. AR/ARMA Transformers uses $L_{I}=512$ for $L_{P}=1$ and $L_{I}=4096$ for $L_{P}=720$. Baselines are consistently set to their best-performing $L_{I}=512$ configuration. Test set MSE and MAE for each model on each setup are reported.
@@
ETTm1 1 0.048 0.140 0.043 0.130 0.042 0.126 0.041 0.123 0.051 0.138 0.043 0.126 0.053 0.140 0.051 0.140 0.054 0.142 0.052 0.140 0.047 0.138 0.049 0.151 0.043 0.127 0.046 0.133 0.044 0.130 0.044 0.129
720 0.396 0.418 0.391 0.416 0.417 0.435 0.408 0.429 0.396 0.419 0.393 0.418 0.395 0.419 0.391 0.416 0.403 0.423 0.399 0.421 0.420 0.412 0.438 0.440 0.418 0.415 0.408 0.422 0.420 0.416 0.802 0.676
ETTm2 1 0.034 0.096 0.030 0.095 0.032 0.094 0.029 0.092 0.033 0.093 0.030 0.092 0.035 0.095 0.033 0.093 0.033 0.095 0.031 0.093 0.034 0.108 0.034 0.112 0.032 0.079 0.033 0.102 0.031 0.098 0.035 0.118
720 0.329 0.373 0.327 0.371 0.341 0.383 0.338 0.381 0.334 0.373 0.328 0.374 0.327 0.372 0.326 0.371 0.331 0.378 0.327 0.373 0.360 0.381 0.369 0.391 0.351 0.379 0.362 0.383 0.403 0.426 2.641 1.387
Weather 1 0.037 0.046 0.032 0.044 0.032 0.042 0.031 0.041 0.033 0.044 0.031 0.041 0.037 0.049 0.034 0.047 0.036 0.045 0.034 0.044 0.038 0.059 0.038 0.064 0.033 0.046 0.035 0.054 0.034 0.052 0.032 0.046
720 0.299 0.328 0.301 0.332 0.308 0.336 0.305 0.337 0.310 0.334 0.299 0.332 0.296 0.326 0.295 0.325 0.299 0.327 0.297 0.326 0.327 0.342 0.317 0.335 0.311 0.334 0.325 0.340 0.319 0.356 0.404 0.423

### v2

@@ 2.8 Selection of $\phi(\cdot)$ and characteristics of implicit MA weights
Table 2: Summary of main TSF results with forecasting horizons $L_{P}\in\{12,24,48,96\}$ and $L_{I}=512$. See Table 8 for the original results. Averages of test set MSE for each model on each dataset are presented. Average rankings (AvgRank) of each model, along with the count of first-place rankings (#Top1), are also included.
Model Pure AR / WAVE Transformer Baseline
@@
Fig. LABEL:fig:vis_phi_selection shows that LeakyReLU provides a balanced lag weight pattern. Unlike ReLU and Sigmoid, which only output values of the same sign, LeakyReLU offers some relaxation while keeping most values negative. This adds flexibility by enabling the desired negative smoothing effect of the MA term, with occasional positive values to enhance modeling flexibility.
To summarize the WAVE attention process with indirect MA weight generation: First, we compute all $\bm{o}^{\text{AR}}_{t}$ using the selected attention mechanism. Then, we apply token shifting and compute all $\bm{r}_{j}$ for $j\leq N-1$. Next, using $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})$, $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$, and $\bm{r}_{j}$, we calculate $\bm{o}^{\text{MA}}_{t}$ with the efficient method matching AR attention, as illustrated in Fig. 4. Finally, the ARMA output is $\bm{o}_{t}=(\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t})\mathbf{W}_{o}$. A summary of MA computation methods for each attention mechanism is in Table 1.
Computational cost and model performance   The introduction of MA term adds three weight matrices $\mathbf{W}^{\text{MA}}_{\{q,k,v\}}$, increasing parameter size. To ensure fair comparison, we use weight-sharing to match the parameter sizes of ARMA and AR models. Specifically, we share $\mathbf{W}_{q}$ between the AR and MA terms and set $\mathbf{W}_{v}$ to an identity matrix, with minimal impact due to the existance of $\mathbf{W}_{o}$ and the MLP layer (see Eq. (1)). This reduces ARMA’s trainable weights to $\mathbf{W}_{q},\mathbf{W}^{\text{AR}}_{k},\mathbf{W}^{\text{MA}}_{k},\mathbf{W}_{o}$, as shown in Fig. 4. While WAVE attention has the same time complexity in order of magnitude as efficient AR attention, its two-stage structure may increase computational costs on constant level. We compare models with different number of layer in the experiements section to show that ARMA’s improved performance is due to structural enhancements, but not increased complexity.
3 Experiments
We conducted comprehensive experiments on 12 widely-used TSF datasets, including Weather, Solar, Electricity (ECL), ETTs, Traffic, and PEMS. See §B.1 for detailed description of datasets.
Table 3: Summary showing that pure AR/WAVE Transformers effectively utilize extended lookback $L_{I}$, while baselines experience performance degradation. $L_{I}\in\{512,1024,2048,4096\}$ with $L_{P}\in\{12,24,48,96\}$ are evaluated and averaged. Original results can be found in Table 10.
Model Pure AR/WAVE Transformer Baseline
+ARMA +ARMA +ARMA +ARMA +ARMA former Former
Table 4: Summary showing that WAVE Transformers with $m=3$ layers consistently outperform their AR counterparts across a wide range of $m$. The same experimental settings and data presentation method as in Table 2 are used. See Table 9 for the original results.
@@ 3 Experiments
      WAVE       Pure AR       Pure AR       Pure AR       Pure AR       Pure AR       Pure AR       Pure AR       Pure AR
@@
Table 5: Summary of long-term time series forecasting for $L_{P}\in\{96,192,336,720\}$. Pure AR/WAVE Transformers uses $(L_{I},L_{P}):(1024,96),(2048,192),(2048,336),(4096,720)$ and we choose the best results for the baselines from $L_{I}\in\{512,1024,2048,4096\}$ for the 4 $L_{P}$ settings. See Fig. 11 for the original results.
Model Pure AR/WAVE Transformer Baseline
@@
Weather 0.221 0.218 0.218 0.215 0.223 0.216 0.220 0.219 0.220 0.218 0.222 0.232 0.216 0.221 0.233 0.251
Solar 0.198 0.195 0.196 0.192 0.204 0.193 0.198 0.195 0.199 0.195 0.209 0.219 0.206 0.202 0.216 0.212
ETTh1 0.414 0.411 0.415 0.411 0.417 0.408 0.409 0.405 0.414 0.410 0.440 0.454 0.408 0.413 0.422 0.906
ETTh2 0.340 0.339 0.343 0.339 0.342 0.340 0.337 0.332 0.348 0.344 0.354 0.374 0.320 0.330 0.426 0.877
ETTm1 0.347 0.345 0.351 0.348 0.357 0.346 0.348 0.345 0.347 0.344 0.354 0.373 0.345 0.346 0.347 0.735
ETTm2 0.249 0.246 0.247 0.243 0.250 0.245 0.246 0.244 0.245 0.240 0.247 0.265 0.243 0.247 0.252 0.576
Baselines   We built AR Transformers using the five attention mechanisms from Table 1 and added MA terms to create WAVE attention for comparison in TSF tasks. Additionally, we included five recent SOTA baselines: FITS (Xu et al., 2024), iTransformer (Liu et al., 2024a), CATS (Lu et al., 2024), PatchTST (Nie et al., 2022), and DLinear (Zeng et al., 2023). We also used a simple channel-dependent encoder-only Transformer, modified by repeating the last input value (like NLinear) to address distribution shift. This model already surpasses older architectures like Autoformer (Wu et al., 2021) and Informer (Zhou et al., 2021), so we excluded these from our comparison.
In the main experiments, both pure AR and WAVE Transformers use a consistent setup: $m=3$ Transformer layers, 8 heads, and model dimension determined by a empirical method $d=16\sqrt{C}$, where $C$ is the number of series. We evaluate their performance using one-step prediction for each test datapoint, aligned with the baselines. Baseline hyperparameters are set to the reported values from their original papers. For more details on hyperparameters and implementation, see §B.2.
We ran all models on all datasets for the four different $L_{P}$. In the main text, we report the average test set MSE for each model across different $L_{P}$ on each dataset and provide the full results in §B.3.
Short-term TSF results   Table 2 highlights the significant performance gains from introducing MA terms to the AR Transformers. All WAVE attention mechanisms outperform their AR counterparts in both average test MSE and ranking, with linear and standard attention showing the best results.
@@
Long-term TSF results  We evaluated pure AR/WAVE Transformers with varying input lengths ($L_{I}$) for different prediction horizons ($L_{P}$): $(L_{I},L_{P}):(1024,96),(2048,192),(2048,336),(4096,720)$. For the baseline models, we selected the best-performing results across multiple input lengths $L_{I}\in\{512,1024,2048,4096\}$ for each prediction horizon. As shown in Table 5, AR models demonstrated comparable performance to baselines, and the incorporation of the ARMA structure consistently yielded improved results over the AR models across all prediction horizons.
Performance of gated linear attention   WAVE brought the greatest improvement to gated linear attention. In gated AR models, the decay factor helps the AR term focus on important local patterns, but it weakens the ability to capture long-term or stable cyclic patterns. By introducing the MA term, local effects are absorbed, allowing the decay factor to function properly in the AR forgetting mechanism, leading to significant performance gains.
Performance of fixed attention   Fixed attention, which lacks dynamic parameter generation, performs worse than other attention. However, its significant improvement with MA terms shows that WAVE enhances the model’s ability structurally to capture comprehensive sequence patterns.
Performance and complexity   The improvement of adding the MA term comes from its ability to model short-term impacts, allowing the AR term to focus on long-term and cyclic effects, not from increased computational costs. Table 4 shows that, regardless of the number of layers $m$ (1 to 8), pure AR Transformers consistently underperform compared to WAVE Transformers with a fixed $m=3$.
Adaptability to Longer $L_{I}$   Previous baseline models typically use $L_{I}$ between 96 and 720, as longer $L_{I}$ often leads to overfitting to long-term patterns, ignoring more important local effects (Zeng et al., 2023; Nie et al., 2022; Liu et al., 2024a). However, the next-step prediction and varying lookback inputs in pure AR/WAVE Transformers help the model focus on tokens closer to the next step, improving generalization. As shown in Table 3, increasing $L_{I}$ from 512 to 4096 improves pure AR/WAVE performance, demonstrating scalability and the ability to properly leverage long-term effects. Also, the WAVE structure consistently boosts AR model performance across different $L_{I}$.
@@
Visualization analysis   Fig. LABEL:fig:vis_test_loss shows test loss curves for different Pure AR/WAVE attention mechanisms on the Weather and ETTm1 datasets, with WAVE consistently outperforming AR in both convergence speed and final loss. Fig. 10 visualizes attention input sequence, AR weights, $\mathbf{B}$, and $\mathbf{\Theta}$ matrices of a test datapoint on Weather, showing how MA weights decouple local patterns, allowing AR weights to focus on cyclic and long-term patterns. Additional visualizations in Figs. LABEL:fig:vis_real_weights_additional_1LABEL:fig:vis_real_weights_additional_4 reinforce that there are important long-term stable seasonal patterns for AR weights to capture that should not be disrupted by applying forget gates or EMA. This explains why gated linear attention underperforms linear attention in our experiments.
@@
Computational cost   Table 7 compares the computational cost of pure AR/WAVE Transformers with baselines on the ETTm1 dataset. Our tokenization method reduces the token size $N$, keeping pure AR/WAVE models’ computational cost comparable to the baselines. Additionally, parameter sharing ensures the MA term doesn’t increase the number of parameters, and the extra FLOPs from using WAVE are not significant.
4 Conclusion, limitation, and future works
We propose the WAVE attention mechanism, which integrates an MA term into existing AR attention using a novel indirect MA weight generation method. This approach maintains the same time complexity and parameter size while ensuring the validity of the implicit MA weights. Experiments demonstrate that WAVE attention successfully decouples and handles long-term and short-term effects. The WAVE Transformer, enhanced with the MA term, outperforms their AR counterparts and achieves state-of-the-art results, offering consistent improvements in training with minimal added computational cost.
One limitation is that we have not explored combining the channel-independent WAVE Transformer with multivariate forecasting models to improve its handling of inter-series relationships. For future work, WAVE attention could be applied to general sequence modeling tasks beyond TSF. Testing on larger-scale datasets, such as using WAVE Transformers for large-scale NLP pretraining, is another promising direction.
Impact Statement
This paper contributes to the field of Machine Learning by presenting a model that enhances the accuracy and efficiency of time series forecasting. The proposed WAVE approach has valuable applications, including improved decision-making in critical domains like transportation and healthcare. Although the societal impacts of this research are largely positive, it is important to ensure responsible implementation and careful oversight, particularly in sensitive applications, to mitigate any potential risks or negative outcomes.
Appendix A Appendix
Appendix B Related works
B.1 Datasets
Traffic Dataset [^7] http://pems.dot.ca.gov/(Wu et al., 2021)
PEMS Dataset [^8] http://pems.dot.ca.gov/(Li et al., 2017)
B.2 Hyper-parameter settings and implementation details
For the hyper-parameter settings of the pure AR/WAVE Transformer, we use $m=3$ Transformer layers, 8 heads, and set the hidden dimension $d$ based on the number of series $C$, using the empirical formula $d=16\lfloor\sqrt{C}\rfloor$. We use $4d$ as the hidden dimension for the feedforward MLP in the Transformer layer. A dropout rate of 0.1 is applied to both the AR term and MA term. We initialize the weights of all linear layers and embedding layers using the GPT-2 weight initialization method, with a normal distribution and a standard deviation of 0.02. For the output projection layers in the attention and MLP, we additionally scale the standard deviation by a factor of $1/\sqrt{m}$, aligned with the GPT-2 setting. Normalization layer is applied both before the input to the Transformer and after the Transformer output. We experimented with both standard LayerNorm and RMSNorm as the normalization layer, finding no significant performance differences, so we opted for RMSNorm for lower computational cost. For token input projection, we use a linear layer to project the $L_{P}$-dimensional token to a $d$-dimensional input vector. In the output projection, we do not tie the weights between the input and output linear layers. A learnable position embedding that maps the integer labels from 1 to $N$ (the input sequence length) to the corresponding $d$-dimensional position vectors is used. At the beginning of the model, we apply RevIN to input series $\mathbf{S}_{I}$, subtracting the mean and dividing by the standard deviation for each series. Before outputting the final result, we multiply by the standard deviation and add the mean back. All input series are processed independently and in parallel, merging different series dimensions into the batch size for parallel computation. The random seed used in all the experiments is 2024.
All training tasks in this paper can be conducted using a single Nvidia RTX 4090 GPU. The batch size is set to 32. For larger datasets, such as Traffic and PEMS07, we use a batch size of 16 or 8, with 2-step or 4-step gradient accumulation to ensure the effective batch size for parameter updates remains 32. During training, pure AR/WAVE Transformers are trained using the next-step prediction objective with MSE loss. We use the AdamW optimizer with betas=(0.9, 0.95) and weight decay=0.1, following the GPT-2 settings. For a fair comparison, the same optimizer is used for training baseline models. It is important to note that the baseline models trained with this AdamW setup show significantly better TSF performance compared to those trained with the default Adam optimizer settings. As a result, the baseline performance presented in this paper may exceed the results reported in their original papers. Since this study focuses on long-term last token prediction results, we apply an additional weight factor to the training loss for the last token, multiplying it by $N$. However, this weighting only slightly affects performance on smaller datasets with fewer data points, such as ETTs, and has little to no effect on larger datasets. Given the minimal impact of this method, the original next-token MSE loss is sufficient for most datasets, without requiring further modifications.
B.3 Supplementary experiment results
@@
Table 9: Results showing that WAVE Transformers with $m=3$ layers consistently outperform their AR counterparts across a wide range of $m$. Forecasting horizons $L_{P}\in\{12,24,48,96\}$ and $L_{I}=512$ are used. Test set MSE and MAE for each model on each experiment setup are presented.
Model WAVE(m=3) Pure AR(m=1) Pure AR(m=2) Pure AR(m=3) Pure AR(m=4) Pure AR(m=5) Pure AR(m=6) Pure AR(m=7) Pure AR(m=8)
@@
Table 10: Results showing that pure AR/WAVE Transformers effectively utilize extended lookback $L_{I}$, while baselines experience performance degradation. $L_{I}\in\{512,1024,2048,4096\}$ with $L_{P}\in\{12,24,48,96\}$ are evaluated. Test set MSE and MAE for each model on each setup are presented.
@@
Table 11: Results of long-term time series forecasting. Baselines are reported with their best-performing results. Test set MSE and MAE for each model on each setup are reported.
@@
Metrics MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE MSE
Weather 96 0.144 0.142 0.142 0.139 0.161 0.142 0.146 0.143 0.147 0.142 0.149 0.158 0.143 0.149 0.150 0.188
192 0.193 0.190 0.190 0.188 0.191 0.191 0.192 0.194 0.188 0.188 0.189 0.203 0.188 0.190 0.211 0.215
336 0.245 0.242 0.239 0.237 0.238 0.236 0.238 0.238 0.236 0.236 0.237 0.250 0.235 0.240 0.255 0.270
720 0.301 0.299 0.300 0.295 0.300 0.296 0.303 0.302 0.309 0.307 0.311 0.316 0.297 0.306 0.316 0.332
Solar 96 0.196 0.192 0.183 0.180 0.209 0.182 0.194 0.191 0.195 0.187 0.189 0.230 0.182 0.209 0.208 0.201
192 0.192 0.191 0.193 0.185 0.201 0.189 0.198 0.194 0.197 0.192 0.206 0.204 0.214 0.192 0.208 0.209
336 0.193 0.191 0.198 0.197 0.199 0.196 0.191 0.191 0.201 0.196 0.219 0.222 0.216 0.200 0.221 0.221
720 0.210 0.206 0.208 0.206 0.206 0.204 0.208 0.205 0.204 0.205 0.221 0.218 0.213 0.205 0.227 0.218
ETTh1 96 0.357 0.360 0.358 0.361 0.378 0.368 0.360 0.356 0.362 0.359 0.369 0.396 0.365 0.370 0.370 0.986
192 0.393 0.391 0.404 0.398 0.401 0.396 0.391 0.389 0.403 0.395 0.435 0.431 0.404 0.412 0.405 0.814
336 0.418 0.415 0.428 0.424 0.419 0.416 0.42 0.418 0.423 0.419 0.468 0.459 0.423 0.422 0.439 0.883
720 0.487 0.478 0.471 0.462 0.469 0.453 0.463 0.458 0.468 0.466 0.488 0.528 0.441 0.447 0.472 0.941
ETTh2 96 0.266 0.268 0.273 0.275 0.285 0.281 0.267 0.263 0.288 0.276 0.270 0.299 0.259 0.274 0.277 1.303
192 0.336 0.339 0.347 0.336 0.335 0.333 0.335 0.329 0.342 0.338 0.348 0.365 0.315 0.339 0.375 0.939
336 0.371 0.366 0.375 0.373 0.363 0.366 0.365 0.357 0.361 0.363 0.376 0.407 0.339 0.329 0.448 0.551
720 0.385 0.382 0.377 0.371 0.385 0.381 0.379 0.379 0.401 0.398 0.421 0.423 0.365 0.379 0.605 0.714
ETTm1 96 0.305 0.301 0.303 0.296 0.336 0.299 0.305 0.301 0.298 0.296 0.305 0.325 0.282 0.290 0.299 0.686
192 0.332 0.329 0.329 0.327 0.332 0.332 0.332 0.329 0.334 0.328 0.334 0.352 0.326 0.328 0.335 0.636
336 0.355 0.354 0.363 0.362 0.360 0.359 0.358 0.356 0.355 0.354 0.363 0.382 0.358 0.359 0.359 0.791
720 0.395 0.395 0.409 0.406 0.398 0.393 0.395 0.393 0.400 0.396 0.412 0.432 0.414 0.405 0.396 0.825
ETTm2 96 0.177 0.174 0.167 0.162 0.178 0.172 0.178 0.174 0.167 0.166 0.164 0.187 0.158 0.165 0.184 0.481
192 0.220 0.218 0.218 0.215 0.216 0.216 0.217 0.217 0.225 0.213 0.211 0.232 0.211 0.214 0.218 0.434
336 0.256 0.256 0.263 0.259 0.264 0.258 0.260 0.254 0.256 0.253 0.259 0.281 0.261 0.266 0.263 0.461
720 0.341 0.337 0.339 0.337 0.340 0.334 0.330 0.329 0.332 0.328 0.352 0.358 0.340 0.344 0.341 0.928